OpenStudy (anonymous):

please help solve –2(x – 3) ≥ 5 – (x + 3)

OpenStudy (rajee_sam):

first distribute -2 to (x-3)

OpenStudy (rajee_sam):

tell me what you get

OpenStudy (anonymous):

-2x-(-6)?

OpenStudy (rajee_sam):

yes which is -2x + 6

OpenStudy (rajee_sam):

on the right hand side eliminate the parentheses by distributing the negative sign

OpenStudy (rajee_sam):

tell me what you get

OpenStudy (anonymous):

-x+(-3)?

OpenStudy (rajee_sam):

5-x-3

OpenStudy (rajee_sam):

which will be 2-x

OpenStudy (rajee_sam):

now \[-2x + 6 \ge 2-x\]

OpenStudy (rajee_sam):

bring all the like terms together

OpenStudy (rajee_sam):

add 2x to the left hand side and right hand side and subtract 2 from both sides.

OpenStudy (rajee_sam):

\[4 \ge x\]

OpenStudy (rajee_sam):

\[x \le 4\]

OpenStudy (anonymous):

Thank you!